OpenStudy (anonymous):

Add and simplify if possible. x-7 x+5 _____ + _____ 2 2 a.) 2x+12 _________ 2 b.) x-1 c.) 2x-2 ________ 2

OpenStudy (anonymous):

b is the answer

OpenStudy (anonymous):

x+x-7+5/2 = 2x-2/2 = 2(x-1)/2= x-1